OpenStudy (anonymous):

A cylinder has a height of 2 meters and a diameter that is 5 times the measure of the height. Using 3.14 for pi, which of the following can be used to calculate the volume? (3.14)(2m)2(10m) (3.14)(2m)2(5m) (3.14)(10m)2(2m) (3.14)(5m)2(2m)

imqwerty (imqwerty):

the last one

OpenStudy (jasmineetta):

the answer is c

imqwerty (imqwerty):

@king_kong_killer

OpenStudy (jasmineetta):

oh yes, it id D.) because the diameter is 5 x 2 which is 10 and the radius of that is 5 :)

imqwerty (imqwerty):

i ws goin to tell that to you :)

OpenStudy (jasmineetta):

:) im pretty bad at math

OpenStudy (anonymous):

thank you both!!!

imqwerty (imqwerty):

Ur welcome

OpenStudy (anonymous):

\[Volume\ of \ Cylinder\ = \pi r^2h\ unit^3\\ as\ given\ \pi= 3.14,height,\ h=2\ m \ | radius,\ r= \frac{ dia }{ 2 }=\frac{ 5\times h }{ 2 }=\frac{ 5\times 2 }{ 2 }=5\ m\\Volume= 3.14\times (5m)^2\times 2m\\clearly\ option \ LAST \ is \ correct!\]